OpenStudy (anonymous):

13 QUESTIONS LEFT!!!! SOMEBODY PLEASE HELP!!!! The inverse of a modulo 39 is b. What is the inverse of 4a modulo 39 in terms of b?

OpenStudy (anonymous):

Shouldn't you know information on this topic to where you don't have to post every single question?

OpenStudy (anonymous):

My opinion.

OpenStudy (anonymous):

sigh. yes you are probably right. but there were 30 questions. i tried.

OpenStudy (anonymous):

Review the topic, find what you don't know much about. I take 60-100 questions and usually only post 5 or less questions I get stuck on.

OpenStudy (anonymous):

Wow! Yous seem really educated. Can you help me though? Trust me; I've tried.

OpenStudy (anonymous):

*you

OpenStudy (anonymous):

To be honest, I have little knowledge of this topic. I will search it up and see how it works, but if I can't help; I'm sorry.

OpenStudy (anonymous):

At least you are trying!

OpenStudy (anonymous):

I tried substituting numbers and seeing if there was a pattern but i got some numbers

OpenStudy (anonymous):

Ops.

OpenStudy (anonymous):

So x + 4a = 0 mod 39. Substract 4a+4b, which is legal because 4(a+b) = 4*0 = 0 all mod 39 Result: x - 4b = 0 mod 39. or x = 4b mod 39.

OpenStudy (anonymous):

I tried that. The website says it's wrong.

OpenStudy (anonymous):

:( sorry

OpenStudy (anonymous):

i've tried 4b, 16b, and 4

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

Sorry, I have no idea. /: